1.) In a sampling distribution of sample slopes, what is each “building block” in the distribution?

a.) the difference between a sample slope and the population slope

b.) a sample mean

c.) a predicted value using the regression equation

d.) a sample slope

e.) a gently sloping turtle

2.) In inference with regression, what does the p-value represent?

a.) the probability that the population slope is zero

b.) the probability that the population slope is not zero

c.) the probability that the population slope is equal to the sample slope

d.) the probability that the sample slope is real

3.) Everything else remaining constant, the larger the sample size:

a.) the smaller the standard error

b.) the smaller the t-value

c.) the smaller value for r-squared

d.) the smaller the slope

4.) We use GSS2008 data to see if age affects the number of times one visited an art museum in the past year. We find, with a sample size of 1496, that both the slope and the value of the standard error are .009. From this, what can you conclude?

a.) The slope is significant at the p<.001 level

b.) The slope is significant at the p<.01 level

c.) The slope is significant at the p<.05 level

d.) The slope is not statistically significant.

5.) We use GSS2008 data to see if age affects the number of times one visited a zoo in the past year. We find, with a sample size of 1496, that the value of the slope is -.019 and the value of the standard error is .004. From this, what can you conclude?

a.) The slope is significant at the p<.001 level

b.) The slope is significant at the p<.01 level

c.) The slope is significant at the p<.05 level

d.) The slope is not statistically significant.
Answer #1
  1. Sample mean
  2. the probability that the sample slope is real
  3. the smaller the standard error
  4. The slope is not statistically significant. (Since Test statistic, t = b1/Sb1 = 0.009/0.009 = 1 hence P-value = 0.1587 so we fail to reject Ho)
  5. The slope is significant at the p<.001 level (Since Test Statistic, t = b1/Sb1 = -0.019/0.004=4.75 hence P-value < 0.00001)